Multiple Issues x2

Thursday, January 26, 2017 11:26 AM

Problem(s) Encountered:

Normally let dewar fill until steady flow however tonight I waited too long and when moving back tank dropped dolly creating hole/chipped paint.  

This is my fault as I should have been paying more attention.  Went to clean up mess and more paint chipped away, not sure if this causes any issues outside of cosmetics

As mentioned thought issues with telescope from previous day were caused by seeing/weather or improper focusing.  Used the prospero script call focus4k#### could not get rid of coma-like appearance.  Followed instructions from Eric''s email however same results when focusing with B filter.  Results were the same with/without filter.  (Same results with OIII filter too.)  Ran movie on 1k and moved through focus just in case I missed something.  I am unaware of anything else I can do to fix this issue and believe it is something with collimation.  Have multiple request for imaging in queue however can not get usable data with these conditions

Went on to try spectroscopy and dome was off 20 degrees each move.  This is a problem I have encountered multiple times and will move dome to home after turning off auto dome, reset leads in dome encoder box, set dome encoders to home (320).  This normally fixes the problem for the night however that was not the case tonight.  This problem reoccurred

three times, each time I reset as above.  Also noted that when homing for end of night dome was off again.  (the yellow light also appears to be out in computer room)

When taking some images noticed double star/ trails.  I have also had this occur in previous run and knew resetting TCS should solve the issue.  Parked scope and followed

During first reset TCS Link was still down after TCINIT in prospero.  Not sure what caused this but tried check-list again, it worked.

Had to establish pointing with the telescope after reset and as mentioned zenith tilt sensor readings require 4K readout to find star.  I don''t remember having to use 4K in previous runs, it''s not a real problem as star is in image but it is time consuming.

Tried to move telescope to -2h HA and heard loud thump.  Went out to check and as suspected the west mirror cover was closed.  Moved telescope to +4H to see if gravity would open cover but did not.  Climbed up and opened cover.  Tried to move telescope to another target and same issue.  I checked gauge in dome and pressure was normal.  I know Eric warned of this in an email however I thought it was only an issue during closing.  Moved telescope back to +4h HA and opened mirror cover again.  Thought about using rope to keep mirror cover open but considering the focus and all other issues tonight decided against it.  When moving back Zenith heard mirror cover close again, went back out and opened to have cover close in correct order.  Moved telescope to ~2h HA to close however mirror covers closed in wrong order.


Earlier this week, the load cell for HP2 of the primary mirror support system failed, along with the power/data link to the pneumatic/electronics interface box (Sussman box on the north side of the mirror cell).  I replaced the load cell, recompiled the code with new calibration coefficients for the new load cell and brought the system back online.  In doing so, I also ended up having to reset the zero point for the load cell for HP1.  I had a feeling that this may impact collimation, and it does appear that it has.  Right now, things appear to be “good enough” for the majority of spectroscopic observations, and even some imaging programs.  That said, collimation will be revisited, likely some time in the 2/1-9 frame.  I hope to use the Andor CCD to collimate since it provides the best feedback to see real time changes relative to fixture adjustments.  In the meantime, I think it is ok that we limp along as is.  Anyone who is in the queue for observations on this current run, especially for those requesting imaging, should be advised of this.

UPDATE, 02/14: On 02/07, collimation was touched up and is now better than it has been for many months.

Not sure how the “O-50” dolly was dropped.  It all checks out ok thankfully.  Care should be implemented when moving the O-50 as the siphon tube inside is plastic (nylon? fiber glass?).  Also, with the tank under ~15psi of pressure, and filled with LN2, any damage to it has the potential to cause serious damage to equipment or worse, personnel.  The paint chips on the platform are superficial and not a big deal.

The dome continues to be an issue.  What I have found with tests today is that when demanded to moved CCW (descending AZ encoder counts), the display often jumps by large amounts (e.g. 320>290>260>255 instead of 320>319>318>317).  Certainly even one of these jumps and the dome will be lost.  When this has occurred, it is only upon initially starting movement and then seems to correct itself.  It has never occurred for CW movement.  A potential simple workaround is to watch the dome AZ readout and if a large jump is noticed, to immediately switch off auto-dome and reset the dome to the telescope azimuth (if the telescope and dome are still co-aligned).  I really am unsure what else to do with this system.  The failures are so intermittent and touching any portion of the system seems to have an impact (positive or negative) that troubleshooting is extremely iterative (read: frustrating).  Worst case, the dome can be controlled manually.  Not ideal, but it’s a work around.

UPDATE, 02/14: Dome encoder issues seem to be highly temperature-dependent.  When I performed subsequent tests, then worked on collimation on the night of the 7th, I saw no dome encoder issues at all.

TCS testing has shown no issues whatsoever today.  Not sure what else to say about this one.

UPDATE, 02/14: Other than tilt sensors values at zenith changing (now +0.11 & +0.07), I have seen no TCS issues.  In fact, telescope pointing/setting has been reliable over the past 10 days and tracking (with guiding disabled) has been steady and positive.

The west mirror cover subsystem failed.  I replaced the two needle valves on the cylinder and initially this appeared to solve the problem.  However, subsequent testing indicates otherwise.  The problem is limited to the west cover.  Air flow goes up to the telescope where it enters a manifold that feeds both the west and east covers (one T-off for each of the two air lines).  Since the problem does not show itself on the east cover, the only other thing it could be is in the lines to the west cylinder, or the cylinder itself.  Testing the lines, valves and cylinder, I find no leaks.  So I’m left thinking that the problem is either with the cylinder, or with an ice blockage.  I’ve been running N2 through the system all afternoon in hopes of mitigating the problem if it is indeed an ice blockage.  We will also look for a spare cylinder as replacing the cylinder is a relatively trivial job.  In the meantime, it seems like the cover will open and close as commanded, even if sometimes it takes a little effort.  And my hopes are that it will start to work more consistently as the temperatures start to rise again.  I’ve spoken with Jessica, and she is ok with tying off the cover to force it to stay open if there is another failure with it.  Not optimal, but it is a work around.  For tonight, I’m tempted to, and may, leave the covers open before I depart the mountain.

UPDATE, 02/14: The issue with the west mirror cover seems to be extremely temperature-dependent.  No issues with the cover at all over the past 10 or more days.